OpenStudy (anonymous):

A sequence of positive real numbers is defined by •a0 = 1, •an+2 = 2an − an+1, for n = 0, 1, 2, ... . Find a2005.

OpenStudy (jamesj):

you need to give us a1 as well.

OpenStudy (zarkon):

because of "sequence of positive real numbers" \[a_1\] will be forced to be 1

OpenStudy (anonymous):

I believe the 2005th term must still be 1. Looking at the equation and creating the characteristic equation we obtain:\[a_{n+2}=-a_{n+1}+2a_n\iff a_{n+2}+a_{n+1}-2a_n=0\]\[\ \iff \lambda^2+\lambda-2=0\iff \lambda_1=-2,\lambda_2=1\] This tells us a closed form for the sequence is in the form:\[a_n=c_1(-2)^n +c_2(1)^n=c_1(-2)^n+c_2\]where the constants depend on initial conditions. However, because this sequence is make of nothing but positive numbers, we already know that:\[c_1=0\], therefore we obtain:\[a_n=c_2\iff a_0=c_2\iff 1= c_2 \iff a_n=1 \] for all n.